OpenStudy (anonymous):

what does ft3 (cubed) mean?

zepdrix (zepdrix):

\[\huge ft^3\] Like this? The 3 let's you know that it is three dimensional. Imagine you wanted to build a house. So you get some wood, you decide that the length of the house will be 50 ft. You do some more measuring, and decide on a depth of 120 ft. Then you do some more measuring. You talk to city counsel and the building contractor. Everyone takes turns screaming at one another! After a long ordeal you finally settle on a height of 40 ft. If you wanted to know how much SPACE is inside of the house, you're asking about VOLUME. We would multiply the length x depth x height. But we don't only multiply the NUMBERS, we also multiply the UNITS (feet in this case). \[\large 50ft \times 120 ft \times 40ft = (50 \times 120 \times 40) (ft \times ft \times ft)\]\[\large = \quad 240,000 ft^3\]